x + y + z = 9
x + 2y + 3z = 23
x + 5y -3Z = -7

Answer by Fombitz(32388)

1.x + y + z = 9
2.x + 2y + 3z = 23
3.x + 5y -3Z = -7
Subtract 2 from 1,
x+y+z-x-2y-3z = 9-23
4.-y-2z=-14
4.y+2z=14
Subtract 3 from 1,
x+y+z-x-5y-3z= 9-(-7)
5.-4y+4z=16
Now you have two eqs. with two unknowns.
4.y+2z=14
5.-4y+4z=16
Multiply eq. 4 by (4)and add to eq. 5 to get rid of y.
4y+8z-4y+4z=14+16
12z=72
z=6
.
.
Then go back and substitute.
y+2z=14
y+12=14
y=2
.
.
x + y + z = 9
x+2+6=9
x=1
(x,y,z)=(1,2,6)